OpenStudy (anonymous):

Solve the equation for y. 7x - 3y = 4

OpenStudy (anonymous):

I choose y = 7x - 4, but other options are y = x + this are the other options y = 7x - 4, y = x + ,y = x -

OpenStudy (anonymous):

is there any \[y=-\frac{ 4-7x}{ 3 }\] ?

OpenStudy (anonymous):

put \[y=\frac{ 4-7x }{ -3 }\] u can then solve x

OpenStudy (anonymous):

A. y = x + B. y = 7x - 4 C. y = x + D. y = x -

OpenStudy (anonymous):

its b guys right...

OpenStudy (anonymous):

it's the closest

OpenStudy (ajprincess):